# balnet: Pathwise Estimation of Covariate Balancing Propensity Scores

### 27 Mar 2026 v2

@@ 1 Introduction
A primary use of the propensity score is inverse probability weighting (IPW), where outcomes are reweighted to construct treated and control groups that are comparable in terms of their pre-treatment covariates. Under the true (oracle) propensity score, IPW exactly balances covariate means between treatment arms and the overall population. However, even without model misspecification, this exact balancing property does not generally hold for propensity scores estimated using standard maximum likelihood logistic regression, which prioritizes accurate prediction of treatment assignment, rather than covariate balance. Imai and Ratkovic (2014) introduced the notion of covariate balancing propensity scores, propensity score estimators explicitly designed to optimize covariate balance. Causal effects estimated using these balancing weights can be more robust to model misspecification, and more efficient than approaches that estimate propensity scores via maximum likelihood (Ben-Michael et al., 2021; Chattopadhyay et al., 2020). These weights can also improve doubly robust methods, such as AIPW (Robins et al., 1994), by stabilizing the debiasing term (van der Laan et al., 2024; Vermeulen and Vansteelandt, 2015).
@@
This observation motivates the use of covariate balancing loss functions with modern regularization techniques. The penalized versions of these loss functions correspond to the primal of optimization problems that constrains covariate imbalance (Ben-Michael et al., 2021; Tan, 2020; Zhao, 2019). As a consequence, the interpretation of the penalty parameter differs fundamentally from that in penalized logistic regression, where it acts as a coefficient budget. For lasso in particular, the penalty parameter directly controls the maximum allowable absolute imbalance across covariates, offering an intuitive metric to diagnose overlap issues, as well as monitor solver progress in applications.
To our knowledge, the closest existing software is RCAL (Tan and Sun, 2020), which implements lasso regularization for covariate balancing loss functions. However, because this solver is written entirely in R, it can be difficult to scale to large datasets. This limitation is particularly salient in applications such as spatial or pixel-level balancing, where both the number of observations and covariates are large.
Our package, balnet for R (R Core Team, 2024), addresses these limitations by combining covariate balancing loss functions with modern optimization techniques for large-scale regularization problems. Our approach uses proximal quasi-Newton methods for generic convex losses, including algorithmic refinements for coordinate descent such as screening and active-set rules (Friedman et al., 2010; Simon et al., 2013; Tibshirani et al., 2012). The core algorithm builds on the adelie C++ elastic net solver by Yang and Hastie (2024), interfaced with R via Rcpp (Eddelbuettel and François, 2011), and relies on numerical linear algebra routines from Eigen (Guennebaud et al., 2010). In an application from Wu et al. (2023) involving spatial balancing with about 140,000 observations and 500 covariates, balnet computes the full regularization path, from raw, unweighted imbalance down to commonly accepted balance thresholds, within minutes on a standard laptop.
@@ 2 Regularization paths for covariate balancing propensity scores
Given suitable overlap conditions, potential outcome means are identified via inverse-propensity weighted (IPW) outcomes. We begin by considering the treated mean, extensions to the average treatment effect (ATE) and the average treatment effect on the treated (ATT) follow by symmetry. The treated mean is identified by
@@
For estimating the ATE, $\mathbb{E}\left[Y_{i}(1)\right]-\mathbb{E}\left[Y_{i}(0)\right]$, balnet fits two propensity score models: one targeting balance for the treated mean and one for the control mean, unlike maximum likelihood which fits one set of coefficients [^2] For software, it is sufficient to implement only the loss in (5). By symmetry, propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ are obtained by solving (5) with inverted treatment indicators $W_{i}^{\prime}=1-W_{i}$. The propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ can also be used to target the ATT, $\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid W_{i}=1\right]$, since the losses only differ by a scaling (the former weights the control to match $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ while the latter weights the controls to match $\frac{1}{\sum_{i=1}^{n}W_{i}}\sum_{i=1}^{n}W_{i}X_{i}$)..
In practice, directly minimizing (5) may not be feasible. Exact balance may be unattainable in settings with high-dimensional covariates ($p>n$) or limited overlap, potentially producing extreme weights. The covariate balancing loss imposes stronger requirements than maximum likelihood estimation, which only requires overlap, e.g., the absence of complete separation.
@@
This condition is stricter than logistic separation: it not only forbids a linear hyperplane from perfectly separating treated and controls, but also prevents the weighted control mean from exceeding the treated mean along any linear combination of covariates. Regularization is therefore key to making this approach practical. We consider the penalized objective
@@
Under lasso penalization, this constraint is relaxed to a box constraint since the corresponding KKT conditions gives
@@
Pathwise optimization is a natural fit for this loss and can provide useful practical diagnostic for overlap. balnet constructs a log-spaced $\lambda$ sequence in the spirit of glmnet, starting at the smallest $\lambda^{\max}$ for which the solution has $\hat{\beta}=0$, and proceeding to a fixed fraction of this value (with default path length 100). For the treated mean, this maximum value is simply given by the largest unweighted treated imbalance,
@@
For lasso penalization, balnet also allows users to specify a target $\lambda_{\min}$ corresponding directly to a maximum allowable imbalance. The path solver then attempts to reach this tolerance and gracefully truncates the path if further reductions are infeasible [^4] If the resulting imbalance remains unsatisfactory, users may consider augmenting the IPW estimator with an outcome model that accounts for the residual imbalance (Athey et al., 2018).. By default, balnet standardizes the covariates, so $\lambda$ can be interpreted as a limit on the standardized mean difference.
@@ 3 Application: Pixel-level balancing
We illustrate balnet using data from Wu et al. (2023), who estimate the causal effect of past low-intensity fire on future high-intensity fires using NASA’s MODIS instrument (Giglio et al., 2003). Treatment status $W_{i}$ is equal to 1 for pixels that experienced a low-intensity fire in a given focal year and 0 otherwise, and the outcome is future high-intensity fire status. To construct a control group, the study conducts a synthetic control analysis (Abadie et al., 2010), using covariate balancing to construct ATT weights such that pre-treatment fire history and geographic covariate trajectories are matched.
@@
fit <- balnet(X, W, target = "ATT", max.imbalance = 0.05,
                verbose = TRUE, num.threads = 4)
Figure 1 summarizes the resulting path fit for a default length of nlambda =100=100555As the path is not automatically truncated before reaching nlambda steps, further reductions in covariate imbalance would be possible if desired.. Motivated by standard diagnostic considerations in IPW analyses (e.g., Austin and Stuart (2015); Greifer (2025); Keele et al. (2025)), some commonly reported metrics are displayed. For a given $\lambda$, the estimated propensity scores targeting the control arm are $\hat{e}^{(0)}(X_{i};\lambda)$ and the ATT weights are
@@
balnet provides extractor methods to seamlessly retrieve these weights for any $\lambda$ on the path. Let $\bar{X}_{\text{target}}$ and $\bar{\sigma}_{\text{target}}$ denote the treated-group covariate means and standard deviations, and let
$\bar{X}_{0}(\lambda)=\frac{\sum_{i:W_{i}=0}\hat{\gamma}_{i}(\lambda)X_{i}}{\sum_{i:W_{i}=0}\hat{\gamma}_{i}(\lambda)}$
denote the weighted control mean. The $p$-vector of standardized mean differences (SMD) is
$\mathrm{SMD}(\lambda)=\frac{\bar{X}_{0}(\lambda)-\bar{X}_{\text{target}}}{\bar{\sigma}_{\text{target}}}.$
@@
Finally, the effective sample size (ESS), expressed as a percentage, is
@@
The results show clear improvements in covariate balance, along with an increased concentration of weights on a smaller subset of control units. The right axis in Figure 1 displays the coefficient of variation of the weights $\hat{\gamma}_{i}(\lambda)$ via the identity linking ESS to the coefficient of variation, $\sqrt{100/\mathrm{ESS}(\lambda)-1}$.
Figure 2 displays the individual SMD using the estimated weights at $\lambda_{\min}$. Figure 2(a) shows the ten covariates with the largest unweighted imbalances, all related to snow water equivalent. Unweighted imbalances are shown as black dots, while weighted imbalances at the selected $\lambda$ are shown in color. By construction of the covariate balancing loss function, the absolute standardized mean difference for each covariate is bounded above by $\lambda$. Visualizing all 500+ covariates in a single panel is impractical, so we supply the groups argument to plot to aggregate covariates by category. Figure 2(b) shows the resulting grouped imbalances.
We conclude by examining the estimated propensity scores at $\lambda_{\min}$ and comparing them with those from regularized logistic regression using maximum likelihood. Figure 3 shows the estimated ATT weights for the control units. The treated fraction is only $\frac{4,466}{141,780}\approx 0.03$, so maximum likelihood logistic regression places most propensity estimates near zero in order to reflect the rarity of treatment, yielding ATT weights with point mass near zero. In contrast, the balancing-loss logistic regression assigns more variable weights so the weighted control group more closely matches the treated group.
Figure 1: Regularization path for covariate balancing using data from Wu et al. (2023) for treatment year 2008 and conifer forests. The plot shows percentage bias reduction (PBR) and effective sample size (ESS) as functions of $\lambda$. The right-hand axis displays the coefficient of variation of the inverse probability weighting (IPW) weights.
@@
Figure 3: Estimated ATT weights using penalized balance loss at $\lambda_{\min}$ and penalized maximum likelihood loss (default $\lambda_{\min}^{\textrm{ML}}$).
@@ 4 Discussion
Regularization is a practical necessity in many modern statistical settings and pathwise solutions are now a standard tool in regularized statistical learning (Friedman et al., 2007, 2010). We use pathwise regularization in a causal setting to target balancing weights for treatment effects under unconfoundedness using logistic covariate balancing loss functions. A natural direction for future work is to extend this approach to additional causal estimands and to panel-data settings where separate balancing weights are estimated along both time and unit dimensions, with regularization paths defined accordingly.
A direction for future algorithmic work is improved tolerance scaling near the end of the regularization path, where full balance may be unattainable and the solver must truncate. Deviance-based scaling, standard in canonical GLM solvers, is not directly applicable here: developing analogous criteria tailored to balancing losses could reduce runtimes. A simple heuristic is to relax the coordinate descent tolerance parameter tol, trading solution accuracy for faster path computation.

### 23 Apr 2026 v3

@@ 1 Introduction
A primary use of the propensity score is inverse probability weighting (IPW), where outcomes are reweighted to construct treated and control groups that are comparable in terms of their pre-treatment covariates. Under the true (oracle) propensity score, IPW exactly balances covariate means between treatment arms and the overall population. However, even without model misspecification, this exact balancing property does not generally hold for propensity scores estimated using standard maximum likelihood logistic regression, which prioritizes accurate prediction of treatment assignment, rather than covariate balance. Imai and Ratkovic (2014) introduced the notion of covariate balancing propensity scores, propensity score estimators explicitly designed to optimize covariate balance. Causal effects estimated using these balancing weights can be more robust to model misspecification and more efficient than approaches that estimate propensity scores via maximum likelihood (Ben-Michael et al., 2021; Chattopadhyay et al., 2020). These weights can also improve doubly robust methods, such as AIPW (Robins et al., 1994), where the IPW weights act as debiasing weights (van der Laan et al., 2024; Vermeulen and Vansteelandt, 2015).
@@
This observation motivates the use of covariate balancing loss functions with modern regularization techniques. The penalized versions of these loss functions correspond to the primal formulation of optimization problems that constrain covariate imbalance (Ben-Michael et al., 2021; Tan, 2020; Zhao, 2019). As a consequence, the interpretation of the penalty parameter differs fundamentally from that in penalized logistic regression, where it acts as a coefficient budget. For lasso in particular, the penalty parameter directly controls the maximum allowable absolute imbalance across covariates, offering an intuitive metric to diagnose overlap issues, as well as monitor solver progress in applications.
To our knowledge, the closest existing software is RCAL (Tan and Sun, 2020), which implements lasso regularization for covariate balancing loss functions. However, because this solver is written entirely in R, it can be difficult to scale to large datasets. This limitation is particularly salient in large-scale applications such as spatial or pixel-level balancing, where both the number of observations and covariates are large.
The package balnet for R (R Core Team, 2024), available at CRAN.R-project.org/package=balnet, is designed to address these limitations by combining covariate balancing loss functions with modern optimization techniques for large-scale regularization problems. Our approach leverages proximal quasi-Newton methods for generic convex losses that incorporates standard algorithmic refinements for coordinate descent such as screening and active-set rules (Friedman et al., 2010; Simon et al., 2013; Tibshirani et al., 2012). The core implementation uses the adelie C++ elastic net solver by Yang and Hastie (2024), interfaced with R via Rcpp (Eddelbuettel and François, 2011), and relies on numerical linear algebra routines from Eigen (Guennebaud et al., 2010). In an application from Wu et al. (2023) involving spatial balancing with about 140,000 observations and 500 covariates, balnet computes the full regularization path, from raw, unweighted imbalance down to commonly accepted balance thresholds, within minutes on a standard laptop.
@@ 2 Regularization paths for covariate balancing propensity scores
Given suitable overlap conditions, potential outcome means are identified via inverse-propensity weighted outcomes. We begin by considering the treated mean, extensions to the average treatment effect (ATE) and the average treatment effect on the treated (ATT) follow by symmetry. The treated mean is identified by
@@
For estimating the ATE, $\mathbb{E}\left[Y_{i}(1)\right]-\mathbb{E}\left[Y_{i}(0)\right]$, there are $2p$ means to balance. Since each logistic model has $p$ free parameters, balnet thus fits two propensity score models: one targeting balance for the treated mean and one for the control mean, unlike maximum likelihood which fits one set of coefficients [^2] It is sufficient to implement only the loss in (5). By symmetry, propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ are obtained by solving (5) with inverted treatment indicators $W_{i}^{\prime}=1-W_{i}$. The propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ can also be used to target the ATT, $\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid W_{i}=1\right]$, since the losses only differ by a scaling (the former weights the control to match $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ while the latter weights the controls to match $\frac{1}{\sum_{i=1}^{n}W_{i}}\sum_{i=1}^{n}W_{i}X_{i}$)..
In practice, directly minimizing (5) may not be feasible. Exact balance may be unattainable in settings with high-dimensional covariates ($p>n$) or under limited overlap, potentially producing extreme weights. The covariate balancing loss imposes stronger requirements than maximum likelihood estimation, which only requires overlap, e.g., the absence of complete separation.
@@
This condition is stricter than logistic separation: it not only rules out a linear hyperplane from perfectly separating treated and controls, but also prevents the weighted control mean from exceeding the treated mean along any linear combination of covariates. Regularization is therefore key to making this approach practical. We consider the penalized objective
@@
Under lasso penalization, this constraint is relaxed to a box constraint since the corresponding KKT conditions give
@@
Pathwise optimization is a natural fit for this loss and can provide a useful practical diagnostic for overlap. balnet constructs a log-spaced $\lambda$ sequence in the spirit of glmnet, starting at the smallest $\lambda^{\max}$ for which the solution has $\hat{\beta}=0$, and proceeding to a fixed fraction of this value (with default path length 100). For the treated mean, this maximum value is simply given by the largest unweighted treated imbalance,
@@
For lasso penalization, balnet also allows users to specify a target $\lambda_{\min}$ corresponding directly to a maximum allowable imbalance. The path solver then attempts to reach this tolerance and gracefully truncates the path if further reductions are infeasible [^4] If the resulting imbalance remains unsatisfactory, users may consider augmenting the weighting estimator with an outcome model to account for the residual imbalance (Athey et al., 2018).. By default, balnet standardizes the covariates, so $\lambda$ directly controls the maximum standardized mean difference.
@@ 3 Application: Pixel-level balancing
We illustrate balnet using data from Wu et al. (2023), who estimate the causal effect of past low-intensity fire on future high-intensity fires using NASA’s MODIS instrument (Giglio et al., 2003). Treatment status $W_{i}$ is equal to 1 for pixels that experienced a low-intensity fire in a given focal year and 0 otherwise, and the outcome is future high-intensity fire status. To construct a control group, the study conducts a synthetic control analysis (Abadie et al., 2010), using covariate balancing to construct ATT weights such that pre-treatment fire history and geographic covariate trajectories match.
@@
R> fit <- balnet(X, W, target = "ATT", max.imbalance = 0.05, verbose = TRUE, num.threads = 4)The verbose option enables a progress bar that reports balance metrics along the $\lambda$ path. The num.threads argument controls the number of cores used by adelie, and balnet for intermediate pre- and post-processing computations, including covariate standardization [^5] Currently, only dense matrix input is supported, and standardization incurs a one-time $n\times p$ copy. The underlying adelie library supports zero-overhead lazy standardization via compile-time specialized matrix types (e.g., sparse and factor representations with optimized access operators and dot product routines (Yang and Hastie, 2024)), enabling memory and computational savings. At the scales considered here, explicit standardization was sufficiently fast. Future work may exploit these specializations to scale to even larger datasets and richer feature representations.. Printing the returned object displays a summary of the path fit with default length nlambda =100=100. As the path is not truncated before reaching nlambda steps, further reductions in covariate imbalance are possible if desired.
R> print(fit)
Call: balnet(X = X, W = W, target = "ATT", max.imbalance = 0.05, verbose = TRUE, num.threads = 4)Control (path: 100/100) Nonzero Avg|SMD| Lambda1 0 2.20052 467.858322 1 2.09474 426.581283 1 1.99770 388.94592...98 57 0.02622 0.0601499 60 0.02415 0.05484100 65 0.02222 0.05000Motivated by standard diagnostics for covariate balance (e.g., Austin and Stuart (2015); Cohn and Zubizarreta (2025); Greifer (2025); Keele et al. (2025)), several commonly reported metrics are displayed. For a given $\lambda$, the estimated propensity scores targeting the control arm are $\hat{e}^{(0)}(X_{i};\lambda)$ and the ATT weights are
@@
Let $\bar{X}_{\text{target}}$ and $\bar{\sigma}_{\text{target}}$ denote the treated-group covariate means and standard deviations, and let
$\bar{X}^{(0)}(\lambda)=\frac{\sum_{i:W_{i}=0}\hat{\gamma}_{i}(\lambda)X_{i}}{\sum_{i:W_{i}=0}\hat{\gamma}_{i}(\lambda)}$
denote the weighted control means. The $p$-vector of standardized mean differences (SMD) is
$\mathrm{SMD}(\lambda)=\frac{\bar{X}^{(0)}(\lambda)-\bar{X}_{\text{target}}}{\bar{\sigma}_{\text{target}}}.$
Along with the number of nonzero coefficients, the print method also shows the average absolute SMD. As a reminder, when using lasso penalization, (10) implies that the maximum absolute standardized imbalance satisfies $|\mathrm{SMD}(\lambda)|_{\infty}\leq\lambda$.
Calling the plot method displays two additional metrics along the $\lambda$ path.
R> plot(fit)
Figure 1: Regularization path for covariate balancing using data from Wu et al. (2023) for treatment year 2008 and conifer forests. The plot shows percentage bias reduction (PBR) and effective sample size (ESS) as functions of $\lambda$. The right-hand axis displays the coefficient of variation of the balancing weights.
@@
The effective sample size (ESS), expressed as a percentage, is
@@
Figure 1 shows clear improvements in covariate balance, along with increased concentration of weights on a smaller subset of control units. The right axis in the figure displays the coefficient of variation of the weights $\hat{\gamma}_{i}(\lambda)$ via the identity $\sqrt{100/\mathrm{ESS}(\lambda)-1}$.
More granular information along the path can be obtained by supplying additional arguments to the plot method. Supplying the lambda argument displays the individual SMDs using the estimated weights at a given $\lambda$. Below, we plot the SMDs for the 10 most imbalanced covariates at $\lambda_{\min}$ (which lambda = 0 interpolates to, as in glmnet).
R> plot(fit, lambda = 0, max = 10)
@@
Figure 2(a) shows the ten covariates with the largest unweighted imbalances, all related to snow water equivalent. Unweighted imbalances are shown as black dots, while weighted imbalances at the selected $\lambda$ are shown in color. By construction, the absolute standardized mean difference for each covariate is bounded above by $\lambda$. Visualizing all 500+ covariates is impractical, so we use the groups argument to plot to aggregate covariates by category. Figure 2(b) shows the resulting grouped imbalances.
R> plot(fit, lambda = 0, groups = covariate.groups)
Finally, in addition to the standard S3 methods for printing, plotting, and predicting, balnet provides an extractor for retrieving the balancing weights for any $\lambda$ on the path. We conclude by displaying the distribution of the estimated balancing weights at $\lambda_{\min}$ in Figure 3.
Figure 3: Estimated balancing ATT weights at $\lambda_{\min}=0.05$.
@@ 4 Discussion
Regularization is a practical necessity in many modern statistical settings and pathwise solutions are now a standard tool in regularized statistical learning (Friedman et al., 2007, 2010). We use pathwise regularization in a causal setting to target balancing weights for treatment effects under unconfoundedness via logistic covariate balancing loss functions.
A natural direction for future work is extensions to additional causal estimands and settings. In longitudinal settings with static treatment, such as panel data, balancing weights are often estimated along both time and unit dimensions, while dynamic treatment settings typically involve sequential optimization. An interesting question is whether pathwise GLM solvers can be adapted to accommodate these settings.
A direction for future algorithmic work is faster handling of regularization path truncation. In settings where exact balance is unattainable, there exists a value of $\lambda$ along the regularization path beyond which the penalized problem diverges. Currently, balnet detects this point when adelie reaches the maximum number of coordinate descent iterations. At the default tolerance level and max iterations, this provides a robust indicator of the path endpoint in our experience, but it requires exhausting the full iteration budget. A simple heuristic is to relax the tolerance (and thus solution precision) and halt when a step produces an explosive increase in the balance loss. Developing a more computationally efficient criteria for early detection of divergence remains for future work.